OpenStudy (anonymous):

Juan is not yet 30 years old. He is 12 years older than Maria. Write an inequality to represent Maria' s age.

OpenStudy (anonymous):

I really need help on this one

OpenStudy (raden):

let J is Juan's age and M is Maria's age first, we have J < 30 and J = M + 12, in other words M = J - 12 supposed the first is an equation J = 30, then subtitute J = 30 to the 2nd equ. M = J - 12 M = 30 - 12 M = 18 therefore, as the conclusion is M < 18 old